OpenStudy (anonymous):

Write 22˚ 15’ 5” using decimal degrees.

OpenStudy (freckles):

\[1'=\frac{1}{60} \deg \\ 1''=\frac{1}{60^2} \deg\]

OpenStudy (freckles):

\[15'=\frac{15}{60} \deg \\ 5''=\frac{5}{60^2} \deg\]

OpenStudy (misty1212):

dinosaur question when was your book written, 1973?

OpenStudy (chrisdbest):

22.251389, 0

OpenStudy (david27):

yea